OpenStudy (anonymous):

Find the area of a 120° sector of a circle whose radius is 6.

jimthompson5910 (jim_thompson5910):

what is the area of the full circle?

OpenStudy (anonymous):

doesnt say

jimthompson5910 (jim_thompson5910):

how would you find it out

jimthompson5910 (jim_thompson5910):

what formula would you use

OpenStudy (anonymous):

i dont know

jimthompson5910 (jim_thompson5910):

Area of a Circle Formula A = pi*r^2 r = radius

jimthompson5910 (jim_thompson5910):

so if r = 6, then what's the area in terms of pi

OpenStudy (anonymous):

\[reqd. area=\pi r ^{2}*\frac{ 120 }{ 360 } \]